# Short-Range Order Based Ultra Fast Large-Scale Modeling of High-Entropy Alloys

### v1

@@ Abstract
High-Entropy Alloys (HEAs) exhibit complex atomic interactions, with short-range order (SRO) playing a critical role in determining their properties. Traditional methods, such as Monte Carlo (MC) and hybrid MC-Molecular Dynamics (MC-MD), are often constrained by system size and computational inefficiency, limiting their applicability for large-scale modeling. We introduce PyHEA, a high-performance Python package with a C++ core, designed to efficiently model HEAs through a combination of global and local search algorithms, incremental SRO calculations, and GPU acceleration. PyHEA delivers an impressive speedup of more than 110,000× over state-of-the-art tools like SCRAPs and ATAT, ensuring accurate SRO calculations with unprecedented efficiency. In practical applications, PyHEA successfully modeled a 256,000-atom Fe-Mn-Cr-Co alloy, replicating literature results and reducing computational time by over tenfold compared to hybrid MC-MD approaches. By eliminating the need for MC-MD hybridization, PyHEA reduces SRO simulation times from hours to minutes, offering a scalable and efficient tool for HEA research. This advancement enables unprecedented scale and efficiency in modeling atomic structures and SRO, accelerating HEA research and development.
Index Terms:
@@
I Introduction
@@
High-Entropy Alloys (HEAs), introduced by Jien-Wei Yeh and colleagues in 2004 1, represent a transformative class in materials science. Comprising five or more principal elements in near-equimolar ratios (typically 5–35 atomic percent each), HEAs exhibit a suite of remarkable properties, including high strength, hardness, and exceptional resistance to oxidation and corrosion 2, 3, 4, 5. These properties make HEAs promising candidates for applications in extreme environments, such as aerospace, nuclear, and energy sectors, while also challenging foundational theories in alloy design 6, 3, 7, 4.
@@
Recent research has revealed the importance of microstructure, especially short-range order (SRO), in determining HEAs’ unique properties 6, 8. particularly short-range order (SRO), which refers to the local atomic arrangement within an alloy 9, 10, 11, 12, 13. SRO, describing the local atomic arrangement, emerges as a prevalent feature in complex concentrated alloys, even under rapid solidification 14. It significantly impacts mechanical performance and thermal stability, enhancing elastic modulus, stiffness, and, in some cases, ductility through an optimized balance of strength and toughness 15, 13, 6, 16, 17. These findings suggest a potential for precise SRO control in tailoring HEA properties, laying a theoretical foundation for new material development.
However, leveraging SRO in HEA design remains a substantial challenge due to the exponentially large configuration space resulting from multiple principal elements 18, 19, 20. For example, in a 250-atom equiatomic Cantor alloy (CrMnFeCoNi), the configuration space is approximately $\frac{250!}{(50!)^{5}}\approx 10^{170}$ — a complexity that escalates with increasing atom count.
Current methods for modeling SRO in HEAs face scalability limitations 21. Simulation tools like ATAT 22, 23 and SCRAPs24, while effective for small to medium-scale systems, encounter efficiency constraints in larger systems and struggle to capture SRO effects, particularly under low-temperature conditions where SRO becomes critical 25, 26, 27. As system size increases, these methods often become computationally prohibitive, limiting their ability to simulate HEAs’ comprehensive behavior at macroscopic scales.
Hybrid Monte Carlo-Molecular Dynamics (MC-MD) methods have been proposed for simulating larger systems of hundreds of thousands of atoms 28, 18, 29. Although these methods support extensive simulations, they demand significant computational resources and rely on highly accurate interatomic potentials, derived from Density Functional Theory (DFT) or Machine Learning Interatomic Potentials (MLIPs) 30, 31, 32, 33. This dependency further exacerbates computational costs, underscoring the need for an efficient modeling framework capable of handling large HEA systems with specific SRO requirements without prohibitive computational demands.
@@
To address these challenges, we introduce PyHEA, a computational framework optimized for large-scale HEA modeling with two core innovations aimed at overcoming scalability and efficiency issues in capturing SRO. First, PyHEA combines global and local search algorithms, which balances broad exploration with detailed refinement, efficiently navigating the expansive configuration space typical in HEAs and significantly reducing modeling times. Second, it leverages high-precision potentials, such as those from DFT and MLIPs, within an incremental SRO calculation framework. This approach allows PyHEA to rapidly evolve SRO values with precision, enabling the construction of accurate, large-scale HEA models within minutes rather than days. By harnessing both robust search techniques and precise local interactions, PyHEA advances HEA modeling capabilities, allowing researchers to conduct simulations at unprecedented scales and detail.
II Results
II-A Parameter Optimization Analysis
@@
This section details the key input parameters and their roles in the optimization process. These parameters include the number of solutions, global iterations, tasks, and search depth, all of which directly influence the balance between computational cost and the quality of the fitness function, defined here as the degree to which the simulated atomic configurations align with target SRO characteristics.
In this context, the fitness function serves as a metric for evaluating the alignment of atomic configurations with the desired SRO properties, where higher fitness values indicate more optimized structures. As illustrated in Fig. 1, the parameters—comprising the number of solutions, global iterations, tasks, and search depth—significantly affect the trade-off between the quality of fitness and computational efficiency.
Parallelism also plays a critical role in the optimization process. As demonstrated in Fig. 1(c), increasing the number of tasks enhances the consistency of fitness outcomes. Nevertheless, execution time grows significantly after 128 tasks, with only marginal improvements in fitness. Therefore, a task count between 64 and 128 represents an optimal range. Similarly, search depth affects the thoroughness of exploration within each iteration. Although deeper searches (greater than 6) yield better fitness outcomes, they come at the expense of increased computational demands. Based on these findings, a search depth of 4 to 6 strikes an optimal balance between precision and computational efficiency.
Consequently, the recommended parameter set for this optimization process comprises 64-128 solutions, 6 global iterations, 64-128 tasks, and a search depth of 4-6. This configuration ensures scalability and computational feasibility while maintaining high precision in fitness assessments.
II-B Comparison with State-of-the-Art Methods
To assess the performance of PyHEA, we conducted a comparative analysis against SCRAPs 24, a state-of-the-art method commonly utilized in HEA modeling. This comparison focused on evaluating both fitness quality and computational efficiency across varying system sizes.
Figs. 2(a) and 2(b) present the fitness distributions for systems with 128 atoms and 2000 atoms, respectively. PyHEA consistently exhibits a more concentrated fitness distribution compared to SCRAPs, particularly as the system size increases. In the case of the 2000-atom system, PyHEA demonstrates a narrower spread with a higher peak frequency around the optimal fitness value, signifying more stable and reliable outcomes. Conversely, SCRAPs shows a broader distribution of fitness values, indicating that PyHEA’s optimization strategy is more robust, especially in larger systems.
@@ II-B Comparison with State-of-the-Art Methods
To further validate PyHEA’s acceleration benefits, we also conducted a comparison with another widely used MC-based tool, ATAT. Table I summarizes the performance metrics across PyHEA (CPU and GPU), SCRAPs, and ATAT. While SCRAPs and ATAT perform comparably for smaller system sizes, they both struggle as the number of atoms increases. Like SCRAPs, ATAT becomes impractical beyond 10,000 atoms. Conversely, PyHEA’s GPU implementation maintains substantially lower execution times across all tested system sizes, underscoring its performance advantages and scalability.
II-C Application
TABLE I: PyHEA computation times (in minutes) for generating optimal structures using different hardware configurations, demonstrating speedups over state-of-the-art methods. The search space consisted of 128 solutions, with the global search running for 10 iterations (typically converging within 3 to 5 iterations). Each iteration employed up to 256 parallel MC tasks, each with a search depth of 10. For simplicity, supercell sizes were set as $S=A\cdot L^{3}$, where L denotes the length of the cubic supercell, and A = 2 for body-centered cubic (bcc) or A = 4 for face-centered cubic (fcc), representing the number of atoms in the primitive cell. Although shells can be included to a range that avoids correlation from periodic boundaries, the short-range order (SRO) was optimized over 3 shells. Comparisons are made with popular MC-based ATAT code23 and the SCRAPs method24. Calculations used a cluster (Intel(R) Xeon(R) Gold 6132 CPU @ 2.60GHz/28-threads) and a NVIDIA 4090 GPU.
  Type Species Atoms ATAT SCRAPs serial [24 cores] PyHEA CPU serial [24 cores] PyHEA GPU CPU speedup over SCRAPs [ATAT] GPU speedup over SCRAPs [ATAT]
  bcc 3 54 $>$144024   2.74[ 0.14]  3.22[0.14] 0.028 0.85[447] 98[51429]
bcc 4 128 $>$10,00024   6.51[ 0.34]  4.19[0.19] 0.028  1.55[2387] 233[357143]
bcc 4 432 -  24.22[ 1.26]  5.62[0.28] 0.029  4.31 835
bcc 4 1024 - 103.14[ 3.18]  6.24[0.32] 0.032 16.53 3223
bcc 4 2000 - 276.00[ 7.02]  7.00[0.36] 0.035 39.43 7886
bcc 5 2000 - 285.48[ 7.27]  9.28[0.48] 0.036 30.11 7930
bcc 5 6750 - 639.45[47.44] 14.37[0.71] 0.046 44.50 13901
bcc 6 27648 - - 39.61[2.05] 0.092 - -
  fcc 3 108 -   17.37[ 0.63]   3.89[ 0.17] 0.005  4.47 3474
fcc 4 256 -  68.77[ 2.24]   5.50[ 0.27] 0.006 12.50 11461
fcc 5 500 - 260.16[ 5.81]   9.39[ 0.43] 0.007 27.71 37166
fcc 5 4000 - 1559.2[64.31]  14.17[ 0.65] 0.014 110.0 111370
fcc 4 256000 - - 1167.2[61.08] 0.577 - -
@@ II-C Application
    Due to excessive computational demands, multinary results cannot be provided.
@@
    The CPU speedup ratio tests is based on the serial results.
@@
The system was simulated in an NPT ensemble for 10 ps, with the Radial Distribution Function (RDF) calculated to analyze the SRO behavior at different temperatures. As depicted in Fig. 3, our results align closely with those from a reference study in which MC simulations were performed on a comparable Fe-Mn-Cr-Co system to investigate compositionally-driven SRO34. Additionally, our simulation outcomes were consistent with the ordering patterns reported in experimental studies 35, 36 which highlights the Fe–Mn/Co/Cr preference in the first coordination shell within an fcc $\text{Fe}_{50}\text{Mn}_{30}\text{Cr}_{10}\text{Co}_{10}$ system.
@@
Employing PyHEA to generate input structures for LAMMPS allowed us to circumvent the extensive MC-MD hybrid computations typically required, which can take several days to complete. PyHEA reduced this computation time to just a few minutes, showcasing a remarkable improvement in computational efficiency. Furthermore, the scale of the simulation, involving more than 200,000 atoms, surpasses the capabilities of most existing methods, emphasizing PyHEA’s robustness and effectiveness in handling large-scale atomistic simulations.
III Discussion
This study demonstrates that PyHEA achieves superior accuracy and can efficiently handle significantly larger system sizes compared to traditional methods. While existing tools such as SCRAPs and ATAT show exponential increases in execution time for systems exceeding 6750 atoms and LAMMPS encounters substantial computational bottlenecks, PyHEA’s GPU acceleration supports simulations with over 200,000 atoms (Table I), delivering performance improvements of up to four orders of magnitude. This scalability is critical for investigating material properties at more realistic scales.
A notable achievement of this work is PyHEA’s ability to substantially reduce the time required for SRO simulations. Traditional approaches utilizing LAMMPS or hybrid MC-MD methods often require several days to complete; in contrast, PyHEA’s optimized workflow reduces simulation time to just a few minutes without sacrificing accuracy (Fig. 5). This level of efficiency not only allows for the simulation of extensive system modeling but also ensures accurate capture of complex atomic interactions and SRO.
In practical applications, PyHEA successfully simulated intricate SRO phenomena in a 256,000-atom Fe-Mn-Cr-Co system across various temperatures. The consistency of PyHEA’s RDF results with those reported in existing literature further supports its accuracy and applicability in real-world scenarios. This underscores PyHEA’s value not only as a computational tool but also as a reliable method for conducting HEA simulations. The ability to conduct simulations at this scale opens the door for future research into more complex, multi-component alloy systems, potentially impacting other fields of materials science and alloy design.
PyHEA’s scalability and performance provide broader implications for the research community. The reduction in simulation time from days to minutes facilitates faster research cycles and reduces the computational resources needed, making large-scale atomistic simulations more feasible and cost-effective, even for teams with limited resources. This capability supports more iterative and exploratory research, where rapid adjustments to simulation parameters and re-analysis can lead to deeper insights and faster progress.
In conclusion, PyHEA represents a significant advancement in HEA modeling by addressing the computational challenges associated with large-scale simulations. By integrating global and local search algorithms, GPU optimization, and incremental SRO calculations, PyHEA surpasses state-of-the-art methods such as SCRAPs, ATAT, and LAMMPS in both accuracy and scalability. Its ability to handle systems with over 200,000 atoms while maintaining computational efficiency establishes it as a leading tool for HEA research. The reduction in simulation time, along with consistent fitness evaluations, facilitates a more comprehensive exploration of material properties, particularly the role of SRO in HEA performance.
Future research could focus on further optimizing PyHEA’s algorithms and expanding their application to larger and more complex material systems. Incorporating advanced machine learning models, such as reinforcement learning, could enhance PyHEA’s adaptive capabilities in optimizing SRO calculations. Additionally, collaborative efforts with experimental teams to validate PyHEA’s predictions under different thermomechanical conditions would strengthen its applicability and predictive accuracy. Ultimately, PyHEA signifies a major step forward in HEA modeling, providing a scalable, accurate, and efficient platform for simulating short-range order and atomic interactions in extensive systems.
IV Methods

### v2

@@ Abstract
High-Entropy Alloys (HEAs) exhibit complex atomic interactions, with short-range order (SRO) playing a critical role in determining their properties. Traditional methods, such as Monte Carlo (MC) based ATAT and SCRAPs, and hybrid Monte Carlo-Molecular Dynamics (MC-MD), are often constrained by system size and computational inefficiency, limiting their applicability for large-scale modeling. We introduce PyHEA, a Python based toolkit with a high-performance C++ core, designed to efficiently model HEAs through a combination of global and local search algorithms, incremental SRO calculations, and GPU acceleration. PyHEA delivers an impressive speedup of more than 333,000× and 10,000× over state-of-the-art tools ATAT and SCRAPs respectively, ensuring accurate SRO calculations with unprecedented efficiency. In practical applications, PyHEA successfully modeled a 256,000-atom Fe-Mn-Cr-Co alloy, replicating literature results and reducing computational time from days to seconds with GPU acceleration, achieving over 1,000× speedup compared to hybrid MC-MD approaches. This huge breakthrough in computational efficiency revolutionizes HEA modeling by bridging theoretical predictions and practical applications, opening new horizons for the design of next-generation HEAs.
keywords:
@@
[…]
[1]organization=School of Software and Microelectronics, Peking University, city=Beijing, country=China
[…]
[2]organization=School of Integrated Circuits, Peking University, city=Beijing, country=China
1 Introduction
@@
High-Entropy Alloys (HEAs), introduced by Jien-Wei Yeh et al. in 2004 [1], have emerged as a transformative class in materials science. Characterized by the presence of five or more principal elements in quasi-equimolar ratios (typically 3 to 55% atomic percent each), HEAs exhibit a unique combination of remarkable properties, including high strength, hardness and exceptional resistance to oxidation and corrosion [2, 3, 4, 5]. These attributes position HEAs as promising materials for applications in extreme environments, such as aerospace, nuclear, and energy sectors. Furthermore, the unconventional nature of HEAs challenges traditional alloy design theories, which spurred significant research interest in their underlying principles [6, 7].
@@
Recent advancements have highlighted the critical role of microstructure in determining HEAs’ extraordinary properties, particularly the influence of short-range order (SRO) [8, 9]. SRO refers to the local atomic arrangement within an alloy, a phenomenon that becomes increasingly prominent in complex concentrated alloys, even under rapid solidification conditions [11, 13]. The impact of SRO on mechanical performance and thermal stability is profound, influencing elastic modulus, stiffness, and ductility by optimizing the balance between strength and toughness [14, 15, 16]. These findings underscore the potential for leveraging SRO control to tailor HEA properties, providing a theoretical framework for the development of new high-performance materials.
The primary challenge in designing HEAs with precise SRO control lies in the vast configurational space. For an equiatomic HEA with $N$ total atoms and $M$ different principal elements, the configurational space size $\Omega$ is given by:
$\Omega=\frac{N!}{\left(\frac{N}{M}!\right)^{M}}.$ (1)
This expression reveals the astronomical complexity of the problem—even a modest Cantor alloy (CrMnFeCoNi) with just 250 atoms yields approximately $10^{170}$ possible configurations [17, 18]. Moreover, realistic applications demand simulations of tens or even hundreds of thousands of atoms to fully capture mechanical and structural properties [33], further amplifying this complexity.
Current approaches to HEA modeling face significant limitations in addressing this challenge. Static structure construction methods, such as ATAT [21, 42, 22] and SCRAPs [23], rely on Special Quasirandom Structures (SQS) and hybrid Monte Carlo approaches, respectively, to generate random alloy configurations. While effective for handling relatively small supercells (tens to hundreds of atoms), these methods become computationally prohibitive for larger systems [20] and fail to capture critical SRO effects, particularly at lower temperatures where atomic ordering significantly influences material properties [24, 25, 26].
Hybrid Monte Carlo-Molecular Dynamics (MC-MD) methods offer an alternative approach by incorporating MC steps into extended MD simulations. Although hybrid MC-MD can theoretically handle systems with hundreds of thousands of atoms [27, 17, 28], its practical effectiveness depends heavily on the availability of highly accurate interatomic potentials. Even when such potentials exist, generating configurations that reliably reflect SRO demands substantial computational resources, often requiring hours or days of calculation time.
To address these challenges, we introduce PyHEA, a high-performance computational framework that revolutionizes large-scale HEA modeling through three key innovations:
  • 1.
    Enhanced accuracy and versatility: PyHEA significantly outperforms existing methods like ATAT and SCRAPs in both modeling efficiency and stability. Our benchmarks in Table 1 demonstrate up to 333,000 and 13,900 times speedup compared with ATAT and SCRAPs respectively while maintaining or improving accuracy. Unlike current tools that primarily generate random configurations, PyHEA supports user-defined target SRO values, making it uniquely suitable for modeling temperature-dependent ordering effects and other practical scenarios where specific atomic arrangements are desired.
  • 2.
    Unprecedented scalability and speed: Through advanced GPU acceleration (CUDA) and optimized search algorithms, PyHEA efficiently handles systems with hundreds of thousands of atoms. For example, it successfully modeled a 256,000-atom Fe-Mn-Cr-Co system in under 40 seconds—a task that typically requires hours or days using traditional hybrid MC-MD method. The details of the hybrid MC-MD simulations are elaborated in section 4. This breakthrough in computational efficiency enables the exploration of mechanical properties and structural characteristics at previously inaccessible scales.
  • 3.
    Comprehensive software ecosystem: Built on Python with a high-performance C++ core, PyHEA combines accessibility with computational efficiency. Released under the LGPL-3.0 license, it provides an end-to-end solution including structure generation, SRO optimization, and analysis tools. The modular design facilitates easy integration with existing workflows and enables rapid customization for specific research needs.
In the following sections, we detail PyHEA’s methodology, demonstrate its performance advantages through comprehensive benchmarks, and showcase its practical applications in a large scale HEA system. Our results establish PyHEA as a transformative tool that bridges the gap between theoretical understanding and practical simulation of large-scale HEA systems, opening new possibilities for materials design and optimization.
2 Results
2.1 Parameter Optimization Analysis
@@
In optimizing atomic configurations for HEAs, several key parameters significantly influence both the quality of results and computational efficiency. These parameters include: the number of solutions (which determines the population size for exploring different atomic arrangements), global iterations (controlling how many times the entire optimization process is repeated), number of tasks (affecting the degree of parallelization in the search process), and search depth (determining how thoroughly each local configuration is explored). Each parameter plays a distinct role in the search algorithm detailed in subsection 4.1 and subsection 4.2. Understanding their optimal values is crucial for achieving the best balance between computational efficiency and solution quality.
To evaluate optimization quality, we employ a fitness function that quantifies how well atomic configurations align with the target Short-Range Order (SRO) characteristics. Lower fitness values indicate better alignment with desired SRO properties. Our analysis reveals distinct patterns in how each parameter affects both fitness and computational efficiency.
Parallelism also plays a critical role in the optimization process. As demonstrated in Fig. 1(c), increasing the number of tasks enhances the consistency of fitness outcomes. Nevertheless, execution time grows significantly after 128 tasks, with only marginal improvements in fitness. Therefore, a task count between 64 and 128 represents an optimal range. Similarly, Fig. 1(d) shows search depth affects the thoroughness of exploration within each iteration. Deeper searches (greater than 8) yield better fitness values, they also come at the expense of increased computational demands. Based on these findings, a search depth of 8-10 strikes an optimal balance between precision and computational efficiency.
Consequently, we recommend two parameter configurations optimized for different priorities. The precision-optimized configuration (128 solutions, 10 global iterations, 256 tasks, search depth 10) achieves maximum accuracy in SRO calculations. The performance-optimized configuration (64 solutions, 6 global iterations, 64 tasks, search depth 8) maintains good optimization quality while reducing computation time as much as possible. Users can select either configuration based on their specific requirements for precision versus computational efficiency.
2.2 Comparison with State-of-the-Art Methods
@@
To evaluate the performance of PyHEA, we conducted a comparative analysis against SCRAPs [23], a state-of-the-art method commonly utilized in HEA modeling. This comparison focused on evaluating both fitness quality and computational efficiency across varying system sizes.
Fig. 2(a-b) present the fitness distributions for typical bcc configurations with 128 atoms and 2000 atoms, respectively. PyHEA consistently exhibits a more concentrated fitness distribution compared to SCRAPs, particularly as the system size increases. In the case of the 2000-atom system, PyHEA demonstrates a narrower spread with a higher peak frequency around the optimal fitness value, signifying more stable and reliable outcomes. Conversely, SCRAPs shows a broader distribution of fitness values, indicating that PyHEA’s optimization strategy is more robust, especially in larger systems.
@@
To further validate PyHEA’s acceleration benefits, we also conducted a comparison with another widely used MC-based tool, ATAT. Table 1 summarizes the performance metrics based on serial execution across PyHEA (CPU and GPU), SCRAPs, and ATAT. While SCRAPs and ATAT perform comparably for smaller system sizes, they both struggle as the number of atoms increases. Like SCRAPs, ATAT becomes impractical beyond a few hundred atoms. Conversely, PyHEA’s GPU implementation maintains substantially lower execution times across all tested system sizes, underscoring its performance advantages and scalability. For a comprehensive comparison of parallel CPU performance between these tools, readers are referred to Table 2 in the Appendix.
Table 1: Performance comparison between PyHEA and existing methods (ATAT and SCRAPs) for optimizing atomic structures in high-entropy alloys, focusing on serial execution performance (parallel performance data available in Appendix). The table shows computation times (in minutes) and speedup ratios for both body-centered cubic (bcc) and face-centered cubic (fcc) structures across varying system sizes (#At) and numbers of species (#Sp). Both PyHEA and SCRAPs were executed in serial mode using a single CPU core (Intel Xeon Gold 6132 @ 2.60GHz), with additional GPU (NVIDIA 4090) results shown for PyHEA. The optimizations used 128 candidate solutions over 10 global iterations with 256 parallel MC tasks (search depth 10). Supercell sizes followed $S=A\cdot L^{3}$, where $A=2$ for bcc or $A=4$ for fcc structures. All calculations optimized short-range order (SRO) over 3 coordination shells. Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT in brackets].
  Type #Sp #At ATAT SCRAPs PyHEA PyHEA Speedup
Serial Serial GPU CPU GPU
  bcc 3 54 >1.4ka 2.7 3.2 .03 .9[438] 98[47k]
bcc 4 128 >10ka 6.5 4.2 .03 1.6[2.4k] 233[333k]
bcc 4 432 24.2 5.6 .03 4.3 835
bcc 4 1024 103 6.2 .03 16.5 3.2k
bcc 4 2048 276 7.0 .04 39.4 7.9k
bcc 5 2048 286 9.3 .04 30.1 7.9k
bcc 5 6.8k 639 14.4 .05 44.5 13.9k
bcc 6 28k 39.6 .09
fcc 3 108 17.4 3.9 .01 4.5 3.5k
fcc 4 256 68.8 5.5 .01 12.5 11.5k
fcc 5 500 260 9.4 .01 27.7 37.2k
fcc 5 4k 1.6k 14.2 .01 110 111k
fcc 4 256k 1.2k .58
@@
    Notation: #Sp = number of elemental species.
@@
    Notation: #At = total number of atoms in the system.
  • c
    Speedup ratios: value vs SCRAPs [value vs ATAT in brackets].
  • d
    ATAT data from Ref. [23]. All times in minutes.
2.3 Application
@@
The system was simulated in an NPT ensemble for 10 ps, with the Radial Distribution Function (RDF) calculated to analyze the SRO behavior at different temperatures. As depicted in Fig. 3, our results align closely with those from a reference study in which MC simulations were performed on a comparable Fe-Mn-Cr-Co system to investigate compositionally-driven SRO [33]. Additionally, our simulation outcomes were consistent with the ordering patterns reported in experimental studies [34, 35] in a $\text{Fe}_{50}\text{Mn}_{30}\text{Cr}_{10}\text{Co}_{10}$ system.
@@
Employing PyHEA to generate input structures for LAMMPS allowed us to circumvent the extensive MC-MD hybrid computations typically required, which can take several days to complete. PyHEA reduced this computation time to just tens of seconds, showcasing a remarkable improvement in computational efficiency. Furthermore, the scale of the simulation, involving 256,000 atoms, surpasses the capabilities of most existing methods, emphasizing PyHEA’s robustness and effectiveness in handling large-scale atomistic simulations.
The temperature-dependent evolution of SRO in this system is further illustrated by Fig. 6 in the Appendix, which presents Warren-Cowley SRO parameter matrices at different temperatures (1373 K, 873 K, and 750 K). These heat maps reveal how the atomic ordering patterns shift from a nearly random distribution at high temperatures (1373 K) to more pronounced local ordering at lower temperatures (750 K), providing valuable insight into the system’s structural evolution.
3 Discussion
This study demonstrates PyHEA’s transformative impact on high-entropy alloy modeling through three key achievements. First, PyHEA achieves unprecedented computational efficiency while maintaining high accuracy. Our benchmarks show that PyHEA outperforms traditional tools like ATAT and SCRAPs by several orders of magnitude, delivering speedups of up to 333,000× and 13,900× respectively (Table 1). This dramatic improvement in performance does not come at the cost of accuracy—PyHEA consistently produces more stable and accurate results across various system sizes, as evidenced by the tighter fitness distributions shown in Fig. 2.
Second, PyHEA successfully addresses the scalability challenge that has long constrained HEA modeling. While existing methods become computationally prohibitive beyond a few thousand atoms, PyHEA’s GPU-accelerated framework efficiently handles systems exceeding 256,000 atoms. This capability represents a significant breakthrough, enabling simulations at experimentally relevant scales that were previously inaccessible. The practical impact of this scalability is demonstrated in our Fe-Mn-Cr-Co system simulation, where PyHEA reduced computation time from days to mere seconds while accurately reproducing temperature-dependent SRO patterns documented in previous studies.
Third, PyHEA’s innovative approach to SRO calculations through incremental updates and GPU optimization establishes a new standard for computational efficiency in materials modeling. The incremental SRO algorithm reduces computational overhead by up to 99.6% (Fig. 5), while GPU acceleration delivers over 1,000× speedup compared to CPU-based calculations. These technical advances make it feasible to conduct large-scale atomistic simulations with unprecedented efficiency, opening new possibilities for materials design and optimization.
The broader implications of PyHEA’s capabilities extend beyond pure computational advancement. The dramatic reduction in simulation time from days to minutes facilitates more iterative and exploratory research approaches. Researchers can now rapidly test hypotheses, adjust parameters, and explore diverse atomic configurations that were previously impractical due to computational constraints. This acceleration of the research cycle has the potential to significantly accelerate the discovery and optimization of new high-entropy alloys.
Furthermore, PyHEA’s success in modeling complex SRO phenomena in large-scale systems provides valuable insights for understanding structure-property relationships in HEAs. The ability to accurately simulate and analyze SRO patterns across different temperatures and compositions, as demonstrated in our Fe-Mn-Cr-Co case study, offers a powerful tool for predicting and optimizing material properties. This capability is particularly crucial given the growing recognition of SRO’s role in determining mechanical and thermal properties of HEAs. Future developments could further enhance PyHEA’s capabilities in several directions. Integration with advanced machine learning models, particularly reinforcement learning approaches, could optimize the search for desired atomic configurations. Expanding the framework to incorporate additional physical properties and interaction models would broaden its applicability across materials science. Additionally, collaborative efforts with experimental teams could strengthen the validation of PyHEA’s predictions under various thermomechanical conditions.
In conclusion, PyHEA represents a significant advancement in computational materials science, bridging the gap between theoretical understanding and practical simulation of large-scale HEA systems. Its combination of unprecedented computational efficiency, accurate SRO modeling, and scalability to experimentally relevant sizes establishes a new paradigm for materials design and optimization. As the field continues to evolve, PyHEA’s framework provides a robust foundation for future innovations in high-entropy alloy research and development.
4 Methods